OpenStudy (anonymous):

A 49kg sample of water absorbs 336kJ of heat.If the water was initially at 25.1Celsius, what is its final temperature?

OpenStudy (jfraser):

same as your last question, just solve for deltaT, not Q

OpenStudy (anonymous):

can you show me how i would perform the question? sorry.

OpenStudy (anonymous):

\[336KG\frac{ ? }{ ? }\]should it be this? >\[\DeltaT= 336KG\]

OpenStudy (anonymous):

SORRY IT WONT LET ME BACK SPACE.

OpenStudy (anonymous):

should the answer be 0.273192943, when rounded 0.27 or 0.3

OpenStudy (anonymous):

49kg x 336kJ x 25.1 C=?

OpenStudy (whpalmer4):

\[Q=mc\Delta{t}\]\[\Delta{t} = \frac{Q}{mc}\] \[\Delta{t} = \frac{336kJ}{49kg*c}\] (I don't recall the value of c, you'll have to supply it) Remember to add \(\Delta{t}\) to the original temperature to get the final temperature.

OpenStudy (whpalmer4):

By the way, look at the units on your proposed multiplication. Does mass*energy*temperature have the right dimensions?